### v1

# WorldSense: Evaluating Real-world Omnimodal Understanding for Multimodal LLMs

Abstract
In this paper, we introduce WorldSense, the first benchmark to assess the multi-modal video understanding, that simultaneously encompasses visual, audio, and text inputs. In contrast to existing benchmarks, our WorldSense has several features: (i) collaboration of omni-modality, we design the evaluation tasks to feature a strong coupling of audio and video, requiring models to effectively utilize the synergistic perception of omni-modality; (ii) diversity of videos and tasks, WorldSense encompasses a diverse collection of 1,662 audio-visual synchronised videos, systematically categorized into 8 primary domains and 67 fine-grained subcategories to cover the broad scenarios, and 3,172 multi-choice QA pairs across 26 distinct tasks to enable the comprehensive evaluation; (iii) high-quality annotations, all the QA pairs are manually labeled by 80 expert annotators with multiple rounds of correction to ensure quality. Based on our WorldSense, we extensively evaluate various state-of-the-art models. The experimental results indicate that existing models face significant challenges in understanding real-world scenarios ( best accuracy). We hope our WorldSense can provide a platform for evaluating the ability in constructing and understanding coherent contexts from omni-modality.
1 Introduction
The ability to comprehend and reason about multimodal inputs—ranging from visual and textual to auditory, tactile, and beyond—is fundamental for both human and artificial agents to navigate and interpret the world. For example, when driving a car, a human driver integrates visual information (e.g., recognizing road signs, traffic lights, and obstacles), auditory cues (e.g., hearing the honking of another car or a siren approaching from behind), and tactile feedback (e.g., the feel of the steering wheel, the vibrations of the road, or the responsiveness of the brakes) to make real-time decisions and ensure safe navigation. This seamless multimodal integration enables intelligent agents to process complex, dynamic environments and respond to subtle cues—an ability that is essential for both human perception and the development of embodied agents designed to interact naturally in the world.
In the recent literature, the development of Multi-modal Large Language Models (MLLMs) [49, 24, 48, 57, 59, 75, 43, 12] have led to remarkable progress on a series of tasks, for example, classification [36], captioning [1, 10, 35], question-answering [56, 50, 40], OCR [45, 76], segmentation [25, 67, 22], autonomous driving [46, 52, 4] and more. However, multi-modal analysis primarily focuses on visual-language information, leaving out crucial modalities like audio, which results in an incomplete evaluation of their multimodal capabilities. While some benchmarks have begun incorporating both visual and audio modalities, they still exhibit several limitations. For example, OmniBench [31] and AV-Odyssey Bench [19] mainly emphasize image evaluation, whereas other benchmarks [18, 29, 70] either restrict to captioning tasks or are limited to simple scenarios, or suffer from low-quality, monotonous questioning patterns.
In this paper, we introduce WorldSense, the first benchmark designed to comprehensively evaluate MLLMs’ ability to perceive, understand, and reason through the integration of omni-modal information in real-world contexts. WorldSense is defined by three key features: (i) Collaboration of Omni-Modality: At the core of WorldSense is the emphasis on the interplay between audio and visual information, as illustrated in Figure 1. Each question is designed to require both modalities for a correct answer; the absence of either modality leads to an incorrect response. Such design rigorously tests a model’s ability to simultaneously process multiple sensory inputs for accurate understanding; (ii) Diversity of Videos and Tasks: WorldSense encompasses 1,662 audio-visual synchronized videos across 8 domains and 67 fine-grained subcategories. It includes 3,172 multiple-choice questions spanning 26 distinct cognitive tasks, from basic perception to complex reasoning. This diversity enables a systematic and comprehensive evaluation of MLLMs’ multimodal understanding capabilities; (iii) High-Quality Annotations: To ensure reliability, all QA pairs in WorldSense are manually labeled by 80 expert annotators and undergo multiple rounds of validation through both human review and automatic MLLM verification. This rigorous process ensures the benchmark’s accuracy and reliability. By integrating these features, WorldSense sets a new standard for evaluating MLLMs’ ability to handle real-world complexity, bridging the gap between artificial intelligence and human-like multimodal understanding.
We conduct comprehensive experiments on WorldSense to evaluate a range of existing MLLMs, including open-source video models, video-audio models, and proprietary models. Our results highlight significant limitations in the current models’ ability to understand and reason through omni-modal information in real-world contexts. First, while open-source video-audio models can process both video and audio inputs, they achieve only around 25% accuracy on our benchmark, comparable to random guessing. In contrast, proprietary models, such as Gemini 1.5 Pro, which processes both video and audio, achieve the highest accuracy at 48%. However, when provided with only a single modality (either video or audio), Gemini 1.5 Pro’s performance drops significantly by approximately 15%, underscoring the crucial role of omni-modal collaboration in real-world tasks. These findings emphasize the strong coupling of modalities in WorldSense and reveal a substantial gap in current MLLMs’ real-world understanding capabilities.
Additionally, we further conduct detailed ablation studies on how different modalities impact model performance. Our analysis shows that visual information plays significant for understanding, while adding audio—whether as raw signals or transcribed subtitles—also improves results. However, raw audio inputs yield superior gains, emphasizing the importance of paralinguistic features such as prosody, intonation, and acoustic context, which are often lost in text transcription. Additionally, increasing the density of video frame sampling, which provides richer temporal information, consistently improves performance. These findings underscore the complementary nature of visual and auditory modalities and highlight the necessity of integrating both for robust real-world understanding.
To summarise, in this paper, we have made three contributions: (i) we introduce, WorldSense, the first benchmark specifically designed to evaluate MLLMs’ omni-modal understanding capabilities in real-world scenarios. WorldSense is characterized by the integration of omni-modality, diverse video categories, and high-quality QA pairs; (ii) We have conducted extensive experiments to evaluate current MLLMs’ capabilities in real-world omni-modal understanding. Our results demonstrate that open-source video-audio MLLMs perform only marginally better than random guessing, while even the best-performing proprietary models achieve only accuracy, revealing significant room for accurate omni-modal understanding; (iii) Through detailed analysis, we investigate the crucial factors affecting real-world omni-modal understanding. Our findings demonstrate that acoustic information, visual cues, and dense temporal sampling all substantially impact model performance. These finds provide valuable guidance for the future development of real-world omni-modal understanding.
2 Related Work
Multimodal Large Language Models. Current Large Language Models (LLMs) are capable of processing multimodal information, including visual, text, and audio. Early works, such as [75, 35, 80, 11, 64, 51], successfully combine vision and text modalities. Subsequent research extends to temporal understanding [66, 24, 58, 39, 61, 65, 26, 14, 69, 74, 60, 6, 41, 34], while parallel efforts [55, 9, 8] focus on audio processing. Recently, researchers shift attention to models [7, 54, 58, 42, 59] capable of simultaneously processing text, vision, and audio inputs. Despite the growing interest in the models which can perform the omnimodality understanding, the absence of a comprehensive evaluation benchmark restricts the development. To address this limitation, we introduce our WorldSense to evaluate models’ capabilities in perceiving and understanding real world omnimodal scenarios.
Multimodal Benchmarks. The development of MLLMs has been driven by benchmarks, evolving from static image understanding [77, 37, 28, 27, 16, 73] to temporal comprehension [30, 38, 53, 79, 13, 17, 23, 63, 68, 72, 33, 3]. However, these benchmarks largely overlook the crucial role of audio in real-world perception. While several audio-visual benchmarks have been proposed, they face significant limitations. AV-Odyssey Bench [19] and OmniBench [31] focus on static images, Music-AVQA [29] and AVQA [70] are domain-specific with monotonous questions, and LongVALE [18] limits its assessment to captioning capabilities alone. Given that existing benchmarks fail to provide a comprehensive evaluation of MLLMs’ real-world understanding capabilities, we introduce WorldSense to address this critical gap in the field.
3 WorldSense
In this section, we aim to detail the construction of WorldSense, including data collection procedure, annotation pipeline, and statistics. Unlike existing benchmarks that assess modalities in isolation, WorldSense evaluates the MLLMs’ ability to perceive, understand, and reason about real-world scenarios through the integration of omni-modal information. As shown in Figure 1, all the multiple-choice questions are carefully crafted, to make sure the questions can only be answered through the comprehensive analysis of text, vision, and audio.
3.1 Design Principle
As for multi-modal evaluation, we base on the audio-visual synchronized videos, which capture temporal events, motion patterns, and audio-visual correlations. To curate the benchmark, we adhere to the following three principles, to ensure a rigorous and comprehensive evaluation.
Benchmarks Modality #Videos Len.(s) #QA Pairs Anno. QA Tokens Sub. Tokens Multi-task Open-domain Sub./Aud. A-V Correlations MSRVTT-QA [68] V 2,990 15.2 72,821 A 8.4 ✗ ✗ ✓ ✗ ✗ ActivityNet-QA [72] V 800 111.4 8,000 M 10.2 ✗ ✗ ✗ ✗ ✗ MVBench [30] V 3,641 16.0 4,000 A 27.3 ✗ ✓ ✓ ✗ ✗ Video-Bench [47] V 5,917 56.0 17,036 A&M 21.3 ✗ ✓ ✓ ✗ ✗ EgoSchema [44] V 5,063 180.0 5,063 A&M 126.8 ✗ ✓ ✗ ✗ ✗ Video-MME [17] V 900 1017.9 2,700 M 35.7 3086.5 ✓ ✓ ✓ ✗ MMBench-Video [13] V 609 165.4 1,998 M 19.3 ✗ ✓ ✓ ✗ ✗ AVQA [70] A+V 57,000 10 57,335 M 14.2 ✗ ✗ ✓ ✓ ✓ Music-AVQA [29] A+V 9,288 60 45,867 M 8.6 ✗ ✗ ✗ ✓ ✓ OmniBench [31] A+I ✗ ✗ 1,142 M 37.8 ✗ ✓ ✓ ✓ ✓ AV-Odyssey [19] A+I ✗ ✗ 4,555 M 19.5 ✗ ✓ ✓ ✓ ✓ LongVALE [18] A+V 8,400 235 ✗ A&M ✗ ✗ ✗ ✓ ✓ ✓ WorldSense A+V 1,662 141.1 3,172 M 37.2 986.2 ✓ ✓ ✓ ✓
Comprehensive Domain. To ensure a thorough evaluation of MLLMs’ real-world understanding, we developed a systematic taxonomy that covers diverse domains and scenarios. This process started with primary categories reflecting core aspects of human experience, which were further subdivided into 67 subcategories to capture specific contexts. This hierarchical structure ensures that our video collection spans a broad range of real-world experiences, providing an ecologically valid basis for assessing multimodal understanding.
Diverse Acoustic Signals. In real-world scenarios, audio signals can be principally categorized into three fundamental types: speech, event and music. Our benchmark incorporates all three to ensure comprehensive coverage, enabling MLLMs to process and understand a wide spectrum of acoustic information—from semantic speech to abstract music and environmental sounds.
Multilevel Assessment. To evaluate the MLLMs’ perceptual and cognitive capabilities, we designed a multi-scale assessment at three levels: recognition (basic audiovisual element detection), understanding (comprehension of multimodal relationships), and reasoning (advanced cognitive tasks like causal inference and abstract thinking). We developed 26 tasks across to assess different aspects of multimodal comprehension, with a focus on integrating audio-visual information at each level.
3.2 Data Collection & Curation
We primarily source our video content from FineVideo [15], a large-scale dataset of high-quality YouTube videos with strong audio-visual correlations across diverse real-world scenarios. To enhance the coverage of musical content, we additionally incorporate videos from MusicAVQA [29].
Our data collection process involves a systematic filtering pipeline to ensure high-quality videos with rich visual-audio semantics and temporal dynamics, as illustrated in Figure 3(a). The process follows three key steps: (i) filtering videos based on predefined taxonomic categories described in Section 3.1 for comprehensive coverage; (ii) utilizing pre-computed metrics, including audio-visual correlation and dynamic content scores, to identify significant clips from an initial pool of approximately 8,000 videos; and (iii) conducting human expert review to assess video quality and real-world relevance. This rigorous process ultimately yields 1,662 high-quality video segments featuring strong audio-visual correlations across diverse real-world contexts.
3.3 Annotation Protocol
QA Annotation. A team of 80 professional annotators is engaged in creating high-quality multiple-choice QA pairs. For each video clip, annotators conduct thorough reviews of both visual and auditory content to ensure comprehensive understanding. They then generate questions and corresponding answer options that specifically require the integration of both visual and audio information for correct responses, thereby enabling effective evaluation of MLLMs’ multimodal understanding capabilities.
Quality Control. To ensure the quality of question-answer pairs, we implement a rigorous quality control process that combines human expertise with automated verification, as illustrated in Figure 3(b). Professional quality control experts evaluate each QA pair based on three essential criteria: (i) linguistic clarity and coherence, (ii) necessity of both visual and audio information for correct answers, and (iii) appropriateness of the question’s difficulty. Questions that fail to meet these standards are returned for revision.
We additionally employ MLLMs for automated verification. Vision-language models like Qwen2-VL[61] verify that questions require multiple modalities for correct answers. Furthermore, multimodal MLLMs capable of processing video, audio, and text, such as Video-LLaMA2[7] and OneLLM [21] are used to assess question difficulty, with questions answered correctly by all models being flagged for manual revision as too simple.
This dual-verification system, combining expert review and automated testing, ensures that all questions in our benchmark are of high-quality and well-formulated, that requires multi-modal comprehension, and present significant challenges for the models.
Methods LLM Size Tech & Science Culture & Politics Daily Life Film & TV Performance Games Sports Music Avg Open-Source Video-Audio MLLMs Unified-IO-2 L [42] 1B 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3 Unified-IO-2 XL [42] 3B 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7 Unified-IO-2 XXL [42] 7B 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9 OneLLM [21] 7B 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8 VideoLLaMA2 [7] 7B 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4 Open-Source Video MLLMs Video-LLaVA [32] 7B 23.6 20.8 19.1 17.3 23.6 17.2 20.8 20.1 20.3 LLaMA3.2 [20] 7B 27.5 25.7 28.9 25.9 27.7 21.1 29.0 26.8 27.1 Qwen2-VL [62] 7B 33.5 29.0 28.4 33.6 30.3 32.3 34.7 38.5 32.4 mPLUG-Owl3 [71] 7B 37.5 31.4 31.0 34.1 33.3 33.2 32.1 30.5 32.9 LLaVA-OneVision [26] 7B 38.9 38.9 36.3 37.6 37.8 37.9 36.3 39.1 37.7 InternVL2.5 [5] 8B 43.7 40.9 34.6 39.7 37.8 36.2 39.4 41.1 39.1 LLaVA-Video [78] 7B 41.6 38.6 40.6 42.1 40.4 39.7 37.0 40.9 40.2 Proprietary MLLMs Claude 3.5 Sonnet [2] - 43.7 31.7 30.6 36.5 30.7 31.9 36.6 33.9 34.8 GPT 4o [24] - 48.0 44.0 38.3 43.5 41.9 41.2 42.6 42.7 42.6 Gemini 1.5 Pro [58] - 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0
3.4 Dataset Statistics
As summarized in Table 1, the WorldSense dataset contains 1,662 video clips with synchronized audio, distributed across 8 primary categories and 67 subcategories. The average duration is 141.1 seconds, with lengths ranging from 30 seconds to over 10 minutes, capturing a wide variety of events and activities. In total, WorldSense includes 3,173 multiple-choice questions, covering three levels.
WorldSense encompasses diverse audio components including speech, environmental sounds, and music. Unlike existing benchmarks that use static images (e.g., AV-Odyssey Bench [19], OmniBench [31]) or feature weak audio-visual correlations (e.g., Video-MME [17]), WorldSense is the first benchmark designed to evaluate MLLMs’ real-world multimodal understanding. It distinguishes itself through: (i) open-domain videos with multi-task evaluation, (ii) original audio-visual content with complete transcriptions, and (iii) carefully crafted questions requiring true audio-visual integration, establishing a more comprehensive benchmark for real-world multimodal understanding assessment.
3.5 Evaluation Paradigm
In our evaluation framework, each test instance consists of a video clip with synchronized audio and a multiple-choice question. Models must process these multi-modal inputs and select the correct answer from several options. Performance is measured by accuracy, comparing the model’s selection to the ground-truth answers. A model’s success is determined by its ability to accurately align with the correct answer.
To rigorously assess the necessity of multimodal integration in real-world understanding, we conduct ablation studies across various modality configurations. This approach not only evaluates overall model performance but also quantifies the models’ reliance on individual modalities, highlighting the critical role of multimodal collaboration in real-world comprehension tasks.
4 Experiments and Findings
This section presents a comprehensive evaluation of both open-source and proprietary MLLMs on the WorldSense benchmark. We first delineate our experimental methodology and evaluation protocols, followed by a comprehensive analysis of quantitative results. Furthermore, we conduct detailed investigations into the important factors that affect performance, providing insights that illuminate potential directions for multi-modal understanding.
4.1 Settings
To comprehensively assess the multi-modal understanding ability, we evaluate three types of MLLMs: (1) open-source audio-visual models, such as Unified-IO-2 [42], OneLLM [21], and VideoLLaMA2 [7]; (ii) open-source MLLMs, such as Qwen2-VL [62], LLaVA-OneVision [26], InternVL2.5 [5], LLaVA-Video [78], and so on; (iii) proprietary MLLMs, such as Claude 3.5 Sonnet [2], GPT 4o [24], and Gemini 1.5 Pro [58]. For all evaluations, we strictly adhere to each model’s official implementation guidelines and the recommended pre-processing procedures. Video frame extraction follows the official configurations specified by corresponding MLLMs, while proprietary models are evaluated according to their API specifications and recommended input formats. Model performance is assessed through direct comparison between model outputs and ground-truth answers.
4.2 Results on WorldSense
Main Results. We present the comprehensive evaluation results on WorldSense in Table 2. Our findings reveal several significant insights regarding the current state of multi-modal models in real-world understanding.
First, current open-source video models are limited in their performance as they process only visual information. This restriction highlights a significant gap in their ability to perform complex, multi-modal understanding tasks, as evidenced by their maximum performance score of only . The results underscore the inadequacies of relying solely on visual processing, emphasizing the need to integrate audio inputs for a more comprehensive and accurate understanding in practical applications.
Methods Modality Tech & Science Culture & Politics Daily Life Film & TV Performance Games Sports Music Avg Unified-IO-2 L [42] Audio 23.0 25.4 24.2 26.7 27.7 23.7 25.0 27.1 25.2 + Caption 21.5 21.1 20.7 17.1 19.9 19.0 22.9 23.7 20.9-4.3 + Video 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3-1.9 Unified-IO-2 XL [42] Audio 21.7 22.4 22.4 22.1 24.7 25.0 25.9 24.7 23.4 + Caption 19.9 19.8 20.8 19.2 20.2 15.9 21.7 25.5 20.7-2.7 + Video 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7+1.3 Unified-IO-2 XXL [42] Audio 27.5 28.7 23.9 23.2 25.8 21.1 26.2 30.2 25.9 + Caption 24.0 26.7 23.0 18.9 18.7 20.7 25.9 29.4 23.7-2.2 + Video 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9+0.0 OneLLM [21] Audio 25.7 26.1 19.3 21.9 25.8 25.9 21.5 22.4 23.0 + Caption 29.6 29.0 25.9 29.1 33.0 26.7 29.2 28.6 28.6+5.6 + Video 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8-0.2 VideoLLaMA2 [7] Audio 23.8 23.4 21.3 22.4 24.7 19.8 27.1 27.9 23.8 + Caption 30.0 30.0 25.6 29.9 28.5 25.0 29.7 29.9 28.5+4.7 + Video 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4+1.6 Gemini 1.5 Pro [58] Audio 40.2 42.9 35.8 33.3 33.0 31.0 33.3 24.7 34.6 + Caption 49.5 52.1 41.8 42.9 46.4 41.8 39.6 36.7 43.6+9.0 + Video 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0+13.4
Second and surprisingly, existing open-source audio-visual MLLMs perform even worse, achieving accuracy rates comparable to random guessing and notably below video-only MLLMs. This counter-intuitive finding reveals that despite having access to both modalities, these models struggle with effective audio-visual integration, suggesting that multimodal processing capability alone does not guarantee better performance without sophisticated integration mechanisms.
Third, among proprietary MLLMs, vision-only models GPT-4o and Claude 3.5 Sonnet perform similarly to the best open-source video MLLMs. Gemini 1.5 Pro, capable of processing both audio and visual information, achieves the highest accuracy of . However, this performance still falls considerably short of requirements for reliable real-world applications, indicating substantial room for improvement.
These comprehensive results illuminate several critical insights: (i) the fundamental importance of audio-visual collaborative understanding in real-world scenarios; (ii) the current significant gap in models’ capabilities for effective multimodal integration, and (iii) the need for more sophisticated approaches to combining and reasoning about multiple modalities. These findings point to crucial directions for future research and development in MLLMs.
Breakdown Results. We conduct a fine-grained analysis of model performance across different audio types and task categories, as shown in Figure 4, revealing important insights into the limitations of existing multimodal models.
First, models consistently underperform on audio-related tasks (e.g., audio recognition, audio counting) compared to other task types, demonstrating significant challenges in audio understanding. Second, spatial reasoning and counting tasks present notable difficulties for current models, a pattern consistently observed across multiple benchmarks. Third, emotion-related tasks prove particularly challenging, likely due to their requirement for integrating subtle and complex multimodal cues, including facial expressions, vocal tones, and contextual speech content. This underperformance in emotional understanding suggests a significant gap in current MLLMs’ training data and capabilities, highlighting an important area for future development.
Additionally, performance varies across audio types. While Gemini 1.5 Pro performs best overall, it shows notably lower accuracy on event-related questions compared to speech or music tasks, possibly due to the complex nature of environmental sounds. Other models also exhibit inconsistent performance across audio types, underscoring a general limitation for existing models for audio understanding.
4.3 Roadmap Towards Real-world Understanding
Given the substantial performance gap revealed in above evaluation, we conduct an in-depth investigation into potential approaches to enhance the MLLMs’ performance.
Vision Information. We investigate the impact of visual information through different input configurations: audio-only, audio with video captions, and audio with video frames. As shown in Table 3, visual information generally improves performance, with Gemini 1.5 Pro’s accuracy increasing from (audio-only) to (+video). However, impact varies across models, with UnifiedIO2 showing inconsistent gains and even degradation with captions.
These findings suggest two important insights: (1) visual information is crucial for enhancing multi-modal understanding when properly integrated, and (2) current models’ ability to effectively utilize visual information remains limited.
Audio Information. We examine the impact of audio information through three configurations: video-only, video with subtitles, and video with original audio.
The results in Table 5 reveal intriguing patterns in how different forms of audio information influence model performance. For Gemini 1.5 Pro, accuracy increases from (video-only) to with subtitles, and further to with original audio. OneLLM shows similar improvements. These results demonstrate that both subtitles and acoustic features (including tone, emotion, and environmental sounds) contribute valuable information for multimodal understanding, beyond what subtitles alone can capture, highlighting the importance of complete acoustic cues for multimodal understanding.
Interestingly, UnifiedIO2 demonstrates performance degradation when integrating either subtitles or audio, with subtitles causing a notable accuracy decline, suggesting difficulties in multimodal processing. Conversely, Video-LLaMA2 improves with both modalities but performs better with subtitles than original audio, indicating a stronger reliance on textual rather than complex acoustic information. We further evaluate video-only MLLMs by providing transcribed subtitles, as shown in Table 5. Nearly all models show significant improvements with subtitle integration, reinforcing the importance of audio information. However, the performance gain is less pronounced in music-related questions, as subtitles cannot effectively capture inherent acoustic features such as melody, rhythm, and harmony.
These evaluations highlight several critical findings: (i) original audio contains rich information beyond what subtitles can capture, particularly for music; (ii) current models show significant limitations in multimodal processing. These insights suggest important directions for improving MLLMs’ ability to integrate acoustic and textual information for comprehensive scene understanding.
Methods Speech Event Music Overall Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Unified-IO-2 L [42] 26.8 13.9-12.9 23.1-3.1 26.9 13.5-13.4 23.2-3.7 26.3 15.0-11.3 24.5-1.8 26.6 14.8-11.8 23.3-3.3 Unified-IO-2 XL [42] 25.0 13.0-12.0 24.6-0.4 24.8 12.3-12.5 25.0+0.2 26.7 15.9-10.8 24.7-2.0 25.3 14.1-11.2 24.7-0.6 Unified-IO-2 XXL [42] 27.0 15.6-11.4 25.7-1.3 26.2 14.2-12.0 25.1-1.1 28.4 19.1-9.3 25.9-2.5 27.2 17.2-10.0 25.9-1.3 OneLLM [21] 12.5 19.6+7.1 23.2+10.7 12.4 19.3+6.9 22.3+9.9 12.4 19.0+6.6 22.5+10.1 12.6 19.6+7.0 22.8+10.2 VideoLLaMA2 [7] 17.1 25.5+8.4 25.2+8.1 16.1 24.9+8.8 24.5+8.4 17.7 27.0+9.3 25.6+7.9 17.4 26.1+8.7 25.4+8.0 Gemini 1.5 Pro [58] 34.3 39.6+5.3 49.2+14.9 33.0 38.9+5.9 48.7+15.7 35.4 39.2+3.8 49.8+14.4 34.4 39.3+4.9 48.0+13.6
Methods Speech Event Music Overall Video + Subtitle Video + Subtitle Video + Subtitle Video + Subtitle Video-LLaVA [32] 20.3 15.4-4.9 19.8 14.4 -5.4 19.5 16.4-3.1 20.3 16.0-4.3 LLaMA3.2 [20] 27.1 29.3+2.2 27.6 29.6+2.0 25.9 28.1+2.2 27.1 28.8+1.7 Qwen2-VL [62] 31.8 41.1+9.3 30.9 39.4+8.5 34.2 41.8+7.6 32.4 41.2+8.8 mPLUG-Owl3 [71] 33.0 39.2+6.2 32.3 38.3+6.0 34.6 39.2+4.6 32.9 38.7+5.8 LLaVA-OneVision [26] 37.7 44.0+6.3 36.3 42.7+6.4 39.7 45.7+6.0 37.7 43.9+6.2 InternVL2.5 [5] 39.0 48.3+9.3 38.6 47.9+9.3 39.2 47.1+7.9 39.1 47.8+8.7 LLaVA-Video [78] 40.5 45.9+5.4 38.9 44.6+5.7 42.3 47.7+5.4 40.2 45.6+5.4 GPT 4o [24] 42.8 51.1+8.3 40.9 50.2+9.3 43.6 49.9+6.3 42.6 50.1+7.5
Video Frames. We investigate the effect of temporal sampling density by varying input frame numbers for video-only MLLMs. As shown in Figure 5, most models demonstrate significant performance improvements with increased frame density, with LLaMA-3.2 being a notable exception. These gains likely result from better capture of fine-grained temporal dynamics and subtle visual changes, emphasizing the importance of dense temporal sampling.
5 Conclusion
In this paper, we introduce WorldSense, the first benchmark designed to evaluate MLLMs’ omnimodal understanding in real-world scenarios. Distinguished by its emphasis on joint omnimodal comprehension across diverse real-world contexts, WorldSense encompasses rich video categories and carefully curated question-answer pairs that necessitate the integration of visual and acoustic information. Through extensive experiments, we expose significant limitations in current MLLMs’ ability to process and coherently integrate omnimodal information. Our analysis demonstrates the importance of omnimodal collaboration in real-world understanding. We hope that WorldSense can serve as a foundational benchmark for advancing human-like omnimodal understanding capabilities.
Appendix A Broader Impacts
Our work on WorldSense has several potential positive impacts on society and AI development, while also presenting certain risks that warrant careful consideration. WorldSense contributes to advancing MLLMs’ ability to understand and interact with the real world through multiple modalities. This progress could benefit various applications, including assistive technologies, educational tools, human-AI interaction systems, safety systems, and so on. We also acknowledge potential risks and challenges. The development of more capable AI systems might raise privacy concerns. Advanced multimodal understanding capabilities could potentially be misused for surveillance or monitoring purposes. We believe that open discussion of these impacts is crucial for responsible development of MLLMs.
Appendix B Ethics Statement
Our research on WorldSense adheres to strict ethical principles and guidelines. We acknowledge several important ethical considerations:
Data Collection and Privacy. All video content in WorldSense has been collected from publicly available sources with appropriate licensing agreements. We have con ducted thorough reviews and implemented comprehensive data processing procedures to ensure privacy protection, including the removal of any personally identifiable information.
Potential Biases. While acknowledging that inherent biases may exist in any dataset, we have undertaken systematic efforts to ensure diverse representation across our video content and question-answer pairs, encompassing various domains, cultures, and contexts. Nevertheless, we recognize that completely eliminating bias remains a significant challenge, and users should carefully consider these potential limitations when utilizing our dataset.
Intended Use. WorldSense is specifically designed to advance research in omnimodal real-world understanding. While we actively encourage the use of this benchmark for academic and research purposes, we strongly caution against any applications that could potentially result in harmful or discriminatory outcomes. Users are expected to adhere to ethical guidelines and responsible practices.
Appendix C License
The WorldSense dataset is released under the CC BY-NC-SA 4.0 License. Authors bear all responsibility in case of violation of rights and confirmation of the data license.